OpenStudy (sherribelle):

A demand function often expresses the price at which x items can be sold. Typically, the lower the price is, the more items there are that can be sold at that price, that is, the greater the demand there is at that price. Suppose p=$400-$0.2x is the demand function for widgets. If we want to sell 100 widgets, where should we set the price? If we can sell 100 widgets for that price, how much total revenue would we receive from the sale of those 100 widgets? Where should we set the price if we want to sell 1200 widgets? If we can sell 1200 widgets for that price, how much total revenue

OpenStudy (faiqraees):

Okay so any thoughts?

OpenStudy (sherribelle):

None, I have troubles setting everything up with word problems.

OpenStudy (sherribelle):

I'm curious to know if widgets = x

OpenStudy (faiqraees):

Yes

OpenStudy (sherribelle):

So, 400-0.2(100) = 400-20 = 380? Then, 400-0.2(1200) = 400-240 = 160?

OpenStudy (sherribelle):

Would you then multiply the answer with x? So 380*100=38,000. And 160*1,200=192,000?

OpenStudy (fortytherapper):

Seems like that would be the process